q=100kL. If the price of capital is r= $120 per day and the price of labor is w= $30 per day, What is the combination of labor and capital if the manufacturer wants to produce 100 units? 


1
Expert's answer
2022-01-19T10:52:29-0500

For optimal combination, MRTS= the ratio of Wage rate to rental rate of capital

"MRTS= \frac{MP_L}{MP_K}=\frac{w}{r}"

"MP_L= \frac{\delta Q}{\delta L}= 100K"


"MP_K= \frac{\delta Q}{\delta K}= 100L"

"\frac{100K}{100L}=\frac{30}{120}\implies L=4K"


Substituting L in the cost production function

100=(100)(K)(4K)

"K^2=1\implies K=1"

L=4K"\implies" L=4